OpenStudy (anonymous):

11*47 in distributive property and multiply?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8

TheSmartOne (thesmartone):

multiple ways to make it distributive property: \(\sf 11*47 = (10+1) *47\) \(\sf 11*47 =11*(50-3)\) are just some of the few...

OpenStudy (anonymous):

thank you @TheSmartOne

TheSmartOne (thesmartone):

np :)